OpenStudy (anonymous):

3^(3x-2) =cube root of 9?

OpenStudy (akashdeepdeb):

Cube root of anything means = anything powered by 1/3 Square root of anything means = anything powered by 1/2 Understanding the above statements? :)

OpenStudy (anonymous):

yesss, what do I do next? :((

OpenStudy (akashdeepdeb):

Okay so how will you write cube root of 9 in terms of the base as 3? :D

OpenStudy (anonymous):

I don't know :( that's where im stuck!

OpenStudy (akashdeepdeb):

You said you understood the above statements. Really? Did you?

OpenStudy (akashdeepdeb):

cube root means powered by 1/3 cube root of 9 is also cube root of 3^2 Or 3^2/3 And on the LHS you are given 3^(3x-2) So 3^(3x-2) = 3^2/3 Bases are = 3 and are same :D So equate the powers 3x - 2 = 2/3 This is just like the preivous question you asked. :) Understood this? :)

OpenStudy (anonymous):

cube root of 9 = 9 to the power of 1/3

OpenStudy (anonymous):

is the answer 0.88 reccurring?

OpenStudy (akashdeepdeb):

EXCELLENT!! :D

OpenStudy (anonymous):

omg im so happy! thanks :)

OpenStudy (akashdeepdeb):

:) Reward medals to those who help you. Go through: http://www.openstudy.com